OpenStudy (anonymous):

@jim_thompson5910 What are all the real zeroes y = (x –12)3 – 7?

jimthompson5910 (jim_thompson5910):

plug in y = 0 and solve for x

OpenStudy (anonymous):

i got 1735=x^3 and then the cube root closest to that is 12^3 which is 1728

jimthompson5910 (jim_thompson5910):

y = (x –12)^3 – 7 0 = (x –12)^3 – 7 7 = (x-12)^3 Keep going until you've isolated x

OpenStudy (anonymous):

i get the same thing... im so confused

jimthompson5910 (jim_thompson5910):

\[\Large 7 = (x-12)^3\] \[\Large \sqrt[3]{7} = x-12\] \[\Large \sqrt[3]{7}+12 = x\] \[\Large x = \sqrt[3]{7}+12 \]

OpenStudy (anonymous):

so then what are the zero's ?

jimthompson5910 (jim_thompson5910):

i just showed you

jimthompson5910 (jim_thompson5910):

that's a zero because if you plug that in for x, you'll get y = 0

OpenStudy (anonymous):

so those are the only two since it will be plus or minus

jimthompson5910 (jim_thompson5910):

no there's only one real root..the other two are imaginary